I got 5 feet per second
"While traveling at a constant speed in a car, the centrifugal acceleration passengers feel while the car is turning is inversely proportional to the radius of the turn. If the passengers feel an acceleration of 4 feet per second per second when the radius of the turn is 100 feet, find the acceleration the passengers feel when the radius of the turn is 200 feet."

Answer by stanbon(75887)

While traveling at a constant speed in a car, the centrifugal acceleration passengers feel while the car is turning is inversely proportional to the radius of the turn.
ca = k/r
-----------------
If the passengers feel an acceleration of 4 feet per second per second when the radius of the turn is 100 feet,
----
4 = k/100
k = 400
-----
Equation:
find the acceleration the passengers feel when the radius of the turn is 200 feet.
----
ca = 400/200
ca = 2 ft/sec
===============
Cheers,
Stan H.
===============